A wire on loading by a weight mg extends in length l. The work done in the process is

Grade:11

1 Answers

Khimraj
6 years ago
Let initial length of wire is Lo and after loqding by a weight mg its length become Lo+l
Initial cetre of mass of wire = Lo/2, final centre of mass = (Lo+l)/2
Work done W = F*(shift in centre of mass)
W = mg*[Lo/2 – (Lo+l)/2]= mgl/2.
Hope it clears. If you like answer then please approve it.